# Forecasting Arctic Temperatures with Temporally Dependent Data Using Quantile Gradient Boosting and Adaptive Conformal Prediction Regions

### 04 Mar 2026 v3

@@ 1 Introduction
A key feature of the IPCC report is its reliance on forecasts at very large spatial and temporal scales, consistent with planet-wide coverage and the gradual pace of climate change. Spatial scales can span thousands of square kilometers. [^1] For example, the Community Earth System Model version 2 (CESM2) is typically run at nominal $1^{\circ}$ resolution, corresponding to grid cells of roughly 10,000–15,000 km2, depending on latitude. Downscaling methods have been proposed for regional grids, but they rest on strong assumptions about the parent Global Climate Model (GCM), and no consensus has emerged on a preferred approach (Nishizawa et al., 2018). Moreover, finer spatial grids would greatly increase computational demands that are already near prohibitive. The temporal scale reflects grid cells updated internally every 30 minutes, with atmosphere–land–ocean–ice coupling on hourly intervals and outputs are usually produced as daily or monthly means (Danabasoglu et al., 2020; National Center for Atmospheric Research, 2020). Projections typically extend years into the future.
The goals of numerical weather prediction (NWP) differ from those of climate modeling and provide output at much finer spatial and temporal detail. [^2] Global weather prediction systems, such as ECMWF’s Integrated Forecasting System (IFS) and NOAA’s Global Forecast System (GFS), use grid cells covering on the order of 80–200 km2. Even the most advanced regional models, such as NOAA’s High-Resolution Rapid Refresh (HRRR) or the Weather Research and Forecasting (WRF) model, employ grid spacings of 1–3 km and still aggregate conditions across areas of 1–9 km2. Their time steps are on the order of seconds to minutes, with outputs every 1–3 hours for global models and hourly (or finer) for regional models (European Centre for Medium-Range Weather Forecasts, 2023; NOAA National Centers for Environmental Prediction, 2023; NOAA Earth System Research Laboratory, 2023). Although NWP calculations are performed at much smaller scales, the resulting forecasts still aggregate across those spatial and temporal units. Site-specific forecasts represent averages across larger areas, and hourly predictions are aggregates over seconds and minutes.
The coarse spatial and temporal resolution characterizing GCMs and NWP is especially consequential for forecasts in high-latitude regions such as the Arctic. Small temperature differences around the local freezing point can have important implications, intensified by “Arctic amplification.” Arctic amplification is discussed in more detail shortly.
In addition, the conditions forecasted are usually standardized. The snowfall expected in, say, Fairbanks, Alaska “during the next 24 hours” is one example. Different locations, however, may wish to supplement such forecasts. In the Arctic, local officials might want to know well in advance the probability that $0^{\circ}\mathrm{C}$ will be exceeded. Such information could guide the timing of sediment removal, preemptive slope closures, and preparation of wastewater systems for meltwater discharge, among other actions.
@@
Estimating forecasting uncertainty is problematic as well (Fu, 2025). Gettleman and Rood (2016, 12) summarize the challenges for climate models, which apply equally to weather forecasting: “Uncertainty in climate models has several components. They are related to the model itself, to the initial conditions of the model, and to the inputs that affect the model. All three must be addressed for the model to be useful.”
There have been recent technical advances (Pathak et al., 2022; Price et al., 2024; Bodnar et al., 2025). Yet issues of coarse spatial scales, standardized output, and equivocal reliability remain largely unresolved (Ortega et al., 2022; Balaji et al., 2022; Chang et al., 2023; Zhang et al., 2025). This paper uses Svalbard, Norway, as a test case to advance Arctic temperature forecasting while addressing these three limitations. Temperatures are the focus because of legitimate concerns about a warming climate.
The statistical methods used here are best viewed as a complement to techniques usually applied at larger spatial and temporal scales. Because these methods are more surgical, they can resolve finer spatial and temporal detail. One can obtain legitimate forecasts coupled with valid measures of forecasting uncertainty well suited for highly local and rapid adaptations to Arctic warming.
@@
Section 2 provides background on Svalbard and the forecasting challenges posed by Arctic amplification. Section 3 describes the data, the construction of holdout samples, and the 14-day lagging of predictors used to support legitimate forecasting. Section 4 introduces the statistical methods, focusing on quantile supervised learning. Section 5 presents the empirical results, emphasizing visualization and algorithm interpretations. Section 6 addresses uncertainty through adaptive conformal prediction regions for multiple time-series data and offers a grounded way to communicate forecast reliability to stakeholders. Section 7 discusses policy implications, with emphasis on short-fuse adaptations to Arctic melting, and Section 8 concludes. A synopsis of the data analysis steps is provided in Appendix A.
@@ 2 Svalbard, Norway: The Forecasting Setting
The maritime climate is distinctive in part because Svalbard is warmed by Atlantic Ocean currents. Average winter temperatures in Longyearbyen range from about $20^{\circ}\mathrm{C}$ to $14^{\circ}\mathrm{C}$, while summers are typically between $3^{\circ}\mathrm{C}$ and $7^{\circ}\mathrm{C}$. In recent years, summer temperatures have more often spiked above $10^{\circ}\mathrm{C}$, a trend climate scientists find troubling.
Svalbard has warmed at a rate several times the global average, with mean temperatures increasing by more than $5^{\circ}\mathrm{C}$ since the 1970s. Glaciers are retreating, sea ice is thinning, and thawing permafrost is reshaping the terrain. The archipelago is now among the world’s regions most affected by climate change (Urbański and Litwicka, 2022; Karlsen et al., 2024; Bradley et al., 2025; Schuler et al., 2025).
2.1 Some Implications for Forecasting Temperatures
A scientific attraction is that Svalbard’s maritime climate presents a demanding test for Arctic temperature forecasting. Unlike continental climates at comparable latitudes, one cannot simply look westward and project that those temperatures will arrive several days later. Immediately to the west lies the Arctic Ocean, not a large landmass with comparable terrain.
Another forecasting challenge is the especially rapid climatic change occurring throughout polar regions such as Svalbard, often termed “Arctic amplification” (Rantanen et al., 2022). The farther one moves from the equator, the faster the rate of warming. Early explanations emphasized the declining albedo caused by sea-ice loss and the increasing exposure of tundra and rock. As more sunlight is absorbed and less reflected, additional heat is retained, reinforcing local warming through a strong positive feedback loop.
Semenov (2021) notes that at least two additional mechanisms are now recognized as important contributors to Arctic amplification. Their most fundamental feature is the Planck effect. In simple terms, the Planck effect describes how efficiently Earth radiates heat into space. That efficiency—the rate at which the outgoing energy flux increases with temperature—is smaller in colder regions. Because the radiative flux follows the Stefan–Boltzmann law, its slope with respect to temperature ($T$) is proportional to $T^{3}$. At the low surface temperatures typical of the polar regions, this slope is relatively flat. As a result, the polar surface radiates infrared energy less effectively than warmer regions experiencing the same temperature increase; heat therefore accumulates more rapidly.
The same Planck effect also operates higher in the atmosphere. The “effective emission height” is the average altitude from which the planet’s infrared radiation escapes directly to space. As greenhouse-gas concentrations rise, this emission level shifts upward. Because temperature generally decreases with altitude, radiation then originates from colder air. By the same $T^{3}$ relationship, the efficiency of infrared emission is reduced. Thus, the Planck effect contributes to polar amplification in two ways: directly at the cold surface and indirectly aloft through the upward shift of the emission height to colder temperatures. Arctic amplification provides a potential complication for the forecasting that follows.
@@ 3 Data
The data come from the Longyearbyen weather station at the airport in Svalbard, Norway, and are easily downloaded from the Integrated Surface Database using the worldmet package in R. The database contains weather station data from around the globe. It uses the same standard format regardless of origin.
Anticipating the use of adaptive conformal prediction regions (Romano, Patterson and Candès, 2019), a variant on split-sample methods is employed. Observations for 2023 constitute the training data. Calibration data from 2022 can provide, in principle, an “honest” assessment of the number of algorithm iterations required (Hyndman and Athanasopoulos, 2021, sec. 5.8). They are also the observations needed to construct proper nonconformal scores.
The calibration data are divided into observations for which the forecasted temperature is $0^{\circ}\mathrm{C}$ and observations for which the forecasted temperature is $\leq 0^{\circ}\mathrm{C}$. The melting of snow, sea ice, glacier ice, and permafrost creates positive feedback loops that change the manner in which temperature variation is produced. On subject-matter and statistical grounds, adaptive conformal methods are therefore applied separately to each subset. [^3] One might wonder why the data splits are made using forecasted temperatures rather than observed temperatures. Observed temperatures are certainly available in the weather station data. But when forecasting, only the forecast is known. If the future temperature were known, there would be no need to forecast it.
Data from 2024 serve as a pristine holdout sample treated as “new cases” to document true forecasting skill. As test data, these observations have no role whatsoever in training. The data split by forecasted temperature applied to the calibration data is applied again. The requirements necessary to capitalize properly on all three datasets within the split-sample approach are discussed in Section 6. Some data from 2021 are used for lagged-variable construction, briefly explained in footnote 4. They play no other role in the analyses.
The training data, calibration data, and test data from the Integrated Surface Database come with hourly observations within each day. Hourly data are too detailed for the analyses to follow and add unnecessary complexity; daily data are used instead.
@@
The response variable is the daily, solar-time 2 p.m. air temperature in degrees Celsius at the Longyearbyen weather station. Each daily temperature value at 2 p.m. solar time is an instantaneous observation—a snapshot of conditions with a constant daily reference time rather than an hourly or daily mean. The 2 p.m. solar-time convention is used uniformly throughout the year, including during the polar night, to provide a consistent temporal reference for forecasting. This differs from global climate and numerical weather prediction procedures, which typically rely on spatially and temporally averaged fields.
@@
Predictors are all lagged by 14 days because that seems to be a lag commonly used by climate scientists (Li et al., 2024). In future work, longer lags could be considered. The weather station predictors include (1) wind direction in degrees from true north, (2) wind speed in meters per second, (3) air temperature in degrees Celsius, (4) atmospheric pressure in hectopascals (hPa), (5) visibility in meters, (6) dew point in degrees Celsius, (7) relative humidity in percent, and (8) a day counter ranging from 1 to 365. The counter can capture temporal trends: on average, the diurnal months are warmer than the nocturnal months, although temperature changes over time may be nonlinear. Several predictors are likely to interact in complex ways (Semenov, 2021). The 14-day lags translate into fitted 2 p.m. temperatures two weeks later that are a foundation for forecasts 14 days in advance. [^4] Lagging variables is a routine procedure in feature construction. In this case, however, lagging the first 14 days of a year pushes their reference points into the final 14 days of the previous year. Data from 2021 are included to provide lagged values for the first 14 days of 2022. Otherwise, the 2021 data are not used.
@@
The complete dataset forms a multiple time series, making temporal dependence a potential complication. Split samples are commonly disjoint subsets chosen at random from the data available, but that is likely to obscure any temporal dependence (Hyndman and Athanasopoulos, 2021, sec. 5.8). A key requirement for the 2022 and 2024 holdout samples is that the same physical processes apply during the identical months of 2022, 2023, and 2024; data for all three years are treated as random realizations from the same underlying joint probability distribution. Nonetheless, the pace of Arctic amplification may affect this comparability, an issue examined empirically in Section 5.2.
@@ 4 Statistical Methods
The analyses to follow produce legitimate forecasts from supervised machine learning applied to temporal data. Holdout samples are used to construct provably valid prediction regions. The statistical framework is an interlocking set of procedures. After a search in Google Scholar, their combination appears to be novel. Each component is discussed in a grounded manner when deployed. Appendix A provides a step-by-step synopsis.
4.1 Background
Machine learning algorithms from statistics and computer science are increasingly used in climate and weather applications, including forecasting (Ma et al., 2023; Miloshevich et al., 2024; Kvånum et al., 2025; Zhang et al., 2025). Especially relevant to the analyses that follow is the work of Velthoen et al. (2023), who analyze several years of daily precipitation observations from the Dutch KNMI weather-station network. The precipitation distribution exhibits a long right tail. Weather-station data provide one source of predictors; deterministic precipitation forecasts from the ECMWF weather model provide another. Combined, these data form a multiple time series.
The authors develop gbex, a quantile gradient-boosting algorithm (Friedman, 2001, 2002) tailored for rare and extreme precipitation events using conditional quantiles at very high levels (e.g., 0.95 or 0.99). Quantile regression methods typically struggle when either tail is sparse. The gbex procedure borrows strength from Extreme-Value Theory (EVT) by fitting a Generalized Pareto Distribution (GPD) to values exceeding a high threshold while allowing the GPD parameters to depend on predictors via quantile gradient boosting. The approach is “validated” by comparing the gbex results with classical methods, such as quantile linear regression (Koenker, 2005), showing superior calibration for the extremes.
There is much to admire in Velthoen et al. (2023), but the analyses to follow depart in several important respects. The setting is above the Arctic Circle, where climate and weather processes differ substantially. Consequently, extreme events are not the primary focus; quantile gradient boosting is employed for other purposes.
No enhancements from parametric statistical procedures are used. As Breiman (2001) emphasizes, there is considerable skepticism about statistical modeling unless it has been thoroughly vetted. Nor is algorithmic tuning employed. Therefore, post–model-selection bias is avoided (Kuchibhotla, Kolassa and Kuffner, 2022).
Time-series data commonly exhibit substantial temporal dependence. To obtain valid assessments of forecasting accuracy and uncertainty, the training procedure is expanded to account for this dependence. The analyses adopt standard time-series methods (Box et al., 2015) as a training extension.
Finally, valid estimates of forecasting uncertainty are obtained using adaptive conformal prediction regions (Romano, Patterson and Candès, 2019). The requisite exchangeability is addressed explicitly. Supplemental results are provided that may be of particular interest to stakeholders.
@@ 4.2 Some Details
With $\tau=0.60$, underestimates receive $0.6/0.4=1.5$ times more weight than overestimates when the loss is computed. For this application, about 40% of the 2023 Longyearbyen temperatures exceed $0^{\circ}\mathrm{C}$. Because relatively high values tend to be underestimated and relatively low values overestimated, fitting the conditional 0.60 quantile forces the gradient-boosting algorithm to work harder to avoid underestimates. This occurs most often among the warmest 40% of temperatures. Indirectly, therefore, these “melting days” are weighted more heavily than the rest. Melting days are a particular concern because practical climate adaptations at a local level are often needed.
If melting days are so important, one might argue for reformulating the analysis as a classification problem—above or below $0^{\circ}\mathrm{C}$. While melting is indeed a critical threshold, the degree to which that threshold is exceeded also matters. The melting process is highly nonlinear and difficult to characterize, but current evidence suggests that melting can increase at an increasing rate as temperatures rise (Polashenski, Perovich and Courville, 2012; Pizner et al., 2024).
@@ 5.1 Univariate Plots
A time series of the 2 p.m. temperatures for 2023 should reflect the expected Arctic seasonal swings. Figure 1 (left panel) shows precisely that. The light-blue irregular line interpolates the daily temperatures, while the smooth black line represents a loess smoother applied to those values. The red horizontal line marks the melting point at $0^{\circ}\mathrm{C}$. Temperatures during the nocturnal months are, on average, colder than during the diurnal months, and transitions between them are gradual. Melting temperatures are common from early June through late September.
@@
The right panel shows a histogram of the 2 p.m. temperatures for 2023 with a generalized extreme value (GEV) distribution overlay. Temperatures range from slightly below $20^{\circ}\mathrm{C}$ to slightly above $10^{\circ}\mathrm{C}$, revealing substantial variability. The GEV overlay adds little interpretive value. The histogram is roughly symmetric, although the left tail is somewhat longer than the right. No clear outliers are apparent.

### 25 Apr 2026 v4

# Forecasting Arctic Temperatures With Quantile Machine Learning

@@ 1 Introduction
Derived from computer simulations such the Community Earth System Model, these are macro-scale conclusions for spatial units from about $10^{5}$ meters to about $10^{8}$ meters. There is a need as well for spatial units at local-scales of $10^{2}$ meters or less (Oke, 1987; Danabasoglu et al., 2020; National Center for Atmospheric Research, 2020). Downscaling methods have been proposed, but they rest on strong assumptions about the parent Global Climate Model, and no consensus has emerged on a preferred approach (Nishizawa et al., 2018).
The most recent numerical weather prediction methods such as NOAA’s Research and Forecasting Model can provide output at local scales (European Centre for Medium-Range Weather Forecasts, 2023; NOAA National Centers for Environmental Prediction, 2023; NOAA Earth System Research Laboratory, 2023), but the grid overlaid takes no account of topographical features or political boundaries. A grid cell might span a coastal mountain range and the Pacific Ocean, averaging over two very different climates. It might also straddle political boundaries — parts of several counties, or portions of two countries — each with different authorities, budgets, and policy instruments. A forecast for such a cell is not merely imprecise; it is operationally useless, because no single decision-making authority has jurisdiction over it.
@@
Estimating forecasting uncertainty also can be problematic as well. Gettleman and Rood (2016, 12) summarize the challenges for climate models that apply equally to weather forecasting: “Uncertainty in climate models has several components. They are related to the model itself, to the initial conditions of the model, and to the inputs that affect the model. All three must be addressed for the model to be useful.” Despite many years of trying, obtaining valid estimates of uncertainty remains problematic (Fu, 2025).
There have been recent advances in climate and weather forecasting (Pathak et al., 2022; Price et al., 2024; Bodnar et al., 2025). Yet issues of coarse spatial scales, and equivocal uncertainty estimates remain largely unresolved (Ortega et al., 2022; Balaji et al., 2022; Chang et al., 2023; Zhang et al., 2025). In this paper, an approach is offered, as a complement to existing climate and weather forecasting, that aims to make some progress on both matters.
Svalbard, Norway, is used as a testbed for which forecasts of Arctic temperatures are made. Temperatures are the focus because of legitimate concerns about a warming climate. The Svalbard data can resolve spatial detail precisely where the forecasts are most needed. Daily forecasts are constructed using quantile supervised learning, that can be seen as a a form of “small AI.” There also are valid measures of forecasting uncertainty well suited for highly local and rapid adaptations to Arctic warming.
@@
Section 2 briefly provides some background on Svalbard. Section 3 describes the data, the construction of holdout samples, and the 14-day lagging of predictors used to support legitimate forecasting. Section 4 introduces the statistical methods, focusing on quantile supervised learning. Section 5 presents the empirical results, emphasizing visualization and algorithm interpretations. Section 6 addresses uncertainty through adaptive conformal prediction regions for multiple time-series data and offers a grounded way to communicate forecast reliability to stakeholders. Section 7 discusses policy implications, with emphasis on short-fuse adaptations to Arctic melting, and Section 8 concludes. A synopsis of the data analysis steps is provided in Appendix A.
@@ 2 Svalbard, Norway: The Forecasting Setting
The maritime climate is shaped substantially by Atlantic Ocean currents. Average winter temperatures in Longyearbyen range from about $20^{\circ}\mathrm{C}$ to $14^{\circ}\mathrm{C}$, while summers are typically between $3^{\circ}\mathrm{C}$ and $7^{\circ}\mathrm{C}$. Temperatures approaching $0^{\circ}\mathrm{C}$ cause melting of sea ice, glaciers, ice sheets and snow cover. Local infrastructure can be threatened in ways that require adaptations of many kinds. Melting permafrost has ecological consequences for plant and animal life. In recent years, summer temperatures have more often spiked above $10^{\circ}\mathrm{C}$.
@@ 3 Data
The data come from the Longyearbyen weather station at the airport in Svalbard, and are easily downloaded from the Integrated Surface Database using the worldmet package in R. The Database contains weather station data from around the globe. It uses the same standard format regardless of origin. The Longyearbyen weather station is meant to inform airport operations and characterize meteorological condition where the vast majority of people live. It less useful beyond a few kilometers outside of Longyearbyen because the terrain changes significantly.
@@
The downloaded data are for the full years of 2022, 2023, and 2024. The response variable is the daily, solar-time 2 p.m. air temperature in degrees Celsius at the Longyearbyen weather station. Each daily temperature value at 2 p.m. solar time is an instantaneous observation—a snapshot of conditions with a constant daily reference time rather than an hourly or daily mean. The 2 p.m. solar-time convention is used uniformly throughout the year to provide a consistent temporal reference for forecasting. It is a good proxy for the highest daily temperature.
@@
The relevant weather station predictors include (1) wind direction in degrees from true north, (2) wind speed in meters per second, (3) air temperature in degrees Celsius, (4) atmospheric pressure in hectopascals (hPa), (5) visibility in meters, (6) dew point in degrees Celsius, (7) relative humidity in percent, and (8) a day counter ranging from 1 to 365. The counter can capture temporal trends: on average, the diurnal months are warmer than the nocturnal months, although temperature changes over time may be nonlinear. Several predictors are likely to interact in complex ways (Semenov, 2021).
Predictors are lagged by 14 days, a choice consistent with some recent climate research (Li et al., 2024) and one that provides operationally useful advance warning for local decision-makers. The 14-day lags translate into fitted 2 p.m. temperatures two weeks later that are a foundation for forecasts 14 days in advance. [^1] Lagging variables is a routine procedure in feature construction. In this case, however, lagging the first 14 days of a year pushes their reference points into the final 14 days of the previous year. Data from 2021 are included to provide lagged values for the first 14 days of 2022. Otherwise, the 2021 data are not used.
Anticipating the use of adaptive conformal prediction regions to properly address forecast uncertainty (Romano, Patterson and Candès, 2019), a variant on split-sample methods is employed. Observations for 2023 constitute the training data. Calibration data from 2022 can provide, under exchangeability, the basis for valid nonconformal scores. The calibration data are divided into observations for which the forecasted temperature is $0^{\circ}\mathrm{C}$ and observations for which the forecasted temperature is $\leq 0^{\circ}\mathrm{C}$. The melting of snow, sea ice, glacier ice, and permafrost creates positive feedback loops that change the manner in which temperature variation is produced. On subject-matter and statistical grounds, adaptive conformal methods are therefore applied separately to each subset. [^2] One might wonder why the data splits are made using forecasted temperatures rather than observed temperatures. Observed temperatures are certainly available in the weather station data. But when real forecasting is, only the forecast is known. If a future temperature were known, there would be no need to forecast it.
Data from 2024 serve as a pristine holdout sample treated as “new cases” to document true forecasting skill. As test data, these observations have no role whatsoever in training. The data split by forecasted temperature applied to the calibration data is applied again. The requirements necessary to capitalize properly on all three datasets within the split-sample approach are discussed in Section 6.
@@
The complete dataset forms a multiple time series, making temporal dependence a potential complication. Split samples are commonly disjoint subsets created by random sampling from the data available, but that is likely to obscure any temporal dependence (Hyndman and Athanasopoulos, 2021, sec. 5.8). A key requirement is that the same physical processes apply during the identical months of 2022, 2023, and 2024; data for all three years are treated as random realizations from the same underlying joint probability distribution. Nonetheless, the pace of Arctic amplification may affect this comparability, an issue examined empirically in Section 5.2.
@@ 4 Statistical Methods
The analyses to follow produce legitimate forecasts from supervised machine learning applied to temporal data. Holdout samples are used for forecasting and to construct provably valid prediction regions. The statistical framework is an interlocking set of procedures. To the author’s knowledge, this combination of methods has not previously been applied to Arctic temperature forecasting. In the next section, each component is discussed in a grounded manner as the procedures are applied. A step-by-step statistical synopsis is provided in the appendix.
@@ 4.1 Some Details
With $\tau=0.60$, underestimates receive $0.6/0.4=1.5$ times more weight than overestimates when the loss is computed. For this application, about 40% of the 2023 Longyearbyen temperatures exceed $0^{\circ}\mathrm{C}$. Because relatively high values tend to be underestimated and relatively low values tend to be overestimated, fitting the conditional 0.60 quantile forces the gradient-boosting algorithm to work harder to avoid underestimates. This occurs most often among the warmest 40% of temperatures. Indirectly, therefore, these “melting days” are weighted more heavily than the rest. Melting days are a particular concern because practical climate adaptations at a local level are often needed.
Adaptive conformal prediction regions are used to obtain, even in finite samples, valid measures of forecast uncertainty. They are adaptive because they capture each forecast’s particular amount of uncertainty. The 2022 calibration data are used with the results the quantile gradient boosting algorithm to obtain residuals that serve as provisional nonconformal scores. They are provisional because their desirable properties depend on the exchangeability; they are only defensible as proper nonconformal scores when exchangeability is not violated for the residuals. [^3] A dataset is exchangeable when the order in the observations are realized does not affect their probability distribution. IID data are exchangeable data, but exchangeable data are not necessarily IID. For the weather station data, temporal dependence in the calibration data residuals, which violates exchangeability, is removed with time series procedures (Box et al., 2015).
The forecast data from 2024 are used to obtain true forecasts based on the results of the quantile gradient boosting algorithm. They are true forecasts because the 2024 data are pristine. They are not used in the training or in the construction of the nonconformal scores. In real forecasting applications, the predictor values are newly realized and therefore unavailable during model training — the same condition that holds here.
@@ 5.1 Univariate Plots
A time series of the 2 p.m. temperatures for 2023 should reflect the expected Arctic seasonal swings. Figure 1 (left panel) shows precisely that. The light-blue irregular line interpolates the daily temperatures, while the smooth black line shows the results of a loess smoother applied to those values. The red horizontal line marks the melting point at $0^{\circ}\mathrm{C}$. Temperatures during the nocturnal months are, on average, colder than during the diurnal months, and transitions between them are gradual. Melting temperatures are common from early June through late September.
@@
The right panel displays a histogram of the 2 p.m. temperatures for 2023 with a generalized extreme value (GEV) distribution overlay. Temperatures range from slightly below $20^{\circ}\mathrm{C}$ to slightly above $10^{\circ}\mathrm{C}$, revealing substantial variability. Although the GEV distribution is frequently applied in studies of high temperatures (McKinnon and Simpson, 2022), it adds little interpretive value here. The histogram is roughly symmetric, although the left tail is somewhat longer than the right. No clear outliers are apparent. The extreme temperatures in both tails are essentially a continuation of the overall distribution. They are not otherwise distinctive.